A large number of neon atoms are in thermal equilibrium. What is the ratio of the number of atoms in a 5s state to number in 3p state at (a) 300K; (b) 600K; (c) 1200K? The energy difference between two states is 1.96 eV. (d) at any of these temperatures, the rate at which a neon gas spontaneously emits 632.8nm radiation is quite low. Explain why?

1
Expert's answer
2021-12-21T12:19:34-0500

a)

"n_1=\frac E{kT_1}=\frac{1.96\cdot 1.6\cdot 10^{-19}}{1.38\cdot 10^{-23}\cdot 300}=75.7,"

b)

"n_2=\frac E{kT_2}=\frac{1.96\cdot 1.6\cdot 10^{-19}}{1.38\cdot 10^{-23}\cdot 600}=37.9,"

c)

"n_3=\frac E{kT_3}=\frac{1.96\cdot 1.6\cdot 10^{-19}}{1.38\cdot 10^{-23}\cdot 1200}=18.9,"

d)

"\frac{hc}{\lambda}=E+\frac{mv^2}2,\implies"

"v=\sqrt{\frac{2(hc-\lambda E)}{\lambda m}}=294~\frac ms ."